The Philae Lander, with a mass of 100 kg, is near to the comet 67P, and has exactly the same velocity. It then fires rockets which supply a force of 20 N for 5.0 s directly away from the comet. (i) What acceleration does it have for this time? (ii) What is the final velocity relative to the comet at the end of 5.0 s.
Answer #1

mass=100 kg

ii. Since the area of accelaration time graph will be = 20*10/100= 1

the change in velocity is 1 m/s

so since earlier they have same speed the relative velocity at end must be 1 m/s

i. accelaration with time is 0.20 m/s2 since a=f/m